OpenStudy (anonymous):

Evaluate each monomial for X = 5, Y = -1, and Z = -4 1. 3X3 2. 2Y2 3.Z2 4. ( YX )2 5. -Yx

OpenStudy (anonymous):

Plzz someone help me on this

OpenStudy (anonymous):

is the question \[3X ^{3}\] or 3X3?

OpenStudy (anonymous):

Its 3X to the 3 power Not 2

OpenStudy (anonymous):

alright, so since x=5 \[3x ^{3}\] = \[3(5)^{3}\] = 3(125) = 375

OpenStudy (anonymous):

Its not 3x2 its 3to the 3rd power.

OpenStudy (anonymous):

\[3^{3}\]?

OpenStudy (anonymous):

Just anwer the Qusetions that i have up there like they are. There POWERS AND PRODUCTS PROBLEMS.

OpenStudy (anonymous):

Well I don't know what you're asking but I'll try my best. For the first one, if it's 3 to the 3rd power the answer is 27

OpenStudy (anonymous):

ok hb the rest

OpenStudy (anonymous):

for number 2 are you asking what 2Y^2 is ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

2Y^2 = 2(-1)^2 = 2(1) = 2

OpenStudy (anonymous):

3. Z^2 = (-4)^2 = 16

OpenStudy (anonymous):

Can you do the rest lol

OpenStudy (anonymous):

4. (YX)^2 = (-1x5)^2 = (-5)^2 = 25

OpenStudy (anonymous):

5. -(Y^x) = -(-1)^5 = -(-1) = 1

OpenStudy (anonymous):

whats Yx

OpenStudy (anonymous):

Y^x = (-1)^5 = -1

OpenStudy (anonymous):

Will thank you for the help now im done with this homework bedtime.